The following sample of nine measurements was randomly selected from a normally distributed population:11,10,8,7,14,9,10,12



Test for significant difference between the sample mean and the population mean 10.Use a=0.05.

1
Expert's answer
2022-05-09T09:33:12-0400
"\mu=\dfrac{11+10+8+7+14+9+10+12}{8}=10.125"


Variance


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=\dfrac{34.875}{7}""s=\sqrt{s^2}=\sqrt{\dfrac{34.875}{7}}\approx2.23207"


1. The following null and alternative hypotheses need to be tested:

"H_0:\mu=10"

"H_a:\mu\not=10"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


2. Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=7" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.364619."


3. The rejection region for this two-tailed test is "R = \{t: |t| > 2.364619\}."


4. The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{10.125-10}{2.23207/\sqrt{8}}\approx0.158397"


5. Since it is observed that "|t| =0.158397<2.364619=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=7" degrees of freedom, "t=0.158397" is "p= 0.878617," and since "p= 0.878617>0.05=\alpha," it is concluded that the null hypothesis is not rejected.


6.Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 10, at the "\alpha = 0.05" significance level.